OpenStudy (anonymous):

simplify 1/2 into -1

OpenStudy (anonymous):

Is this possible? Have you missed out an x?

OpenStudy (anonymous):

well, this expression reads like this, the 1/2 is in a parenthesis and the -1 is an exponent of it outside the paranthesis

OpenStudy (anonymous):

You mean like \[(1/2)^{-1}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok. Well, a negative integer exponent can be written, thus: \[a ^{-n}= 1/a^n\] So, \[(1/2)^{-1} = 1/(1/2)\] which will be 2. Hope this helps you understand it a little better.

OpenStudy (anonymous):

You know, I am not sure of how to type the exponents into this area so I will just write the problem for you. I hope you understand. What about simplifying an X with an exponent of 2 by a X with an exponent of 3. Oh, by the way thank you for the first problem, it really helps a lot. Thank you very much.

OpenStudy (anonymous):

see you later

OpenStudy (anonymous):

\[x^2 / x^3\] If, so, then written out in a longer hand will make it easier to understand: \[(x \times x) /( x \times x \times x)\] so, four of the x will cancel out, simplifying the question to \[1/x\]